OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (johnweldon1993):

You need to first distribute the numbers outside the parenthesis...into the parenthesis...this means you take that number...and multiply it by each term in the parenthesis...so 2(x + 6) − 10 = 12 + 4(x − 1) ( ^ ) is 2 times x = ? 2 times 6 = ? The answer to those 2 questions are... 2x and 12 so now we have 2x + 12 − 10 = 12 + 4(x − 1) ( ^ ) this one is written as 4 times x = ? 4 times -1 = ? the answer to THOSE 2 questions are 4x and -4 so now we have 2x + 12 − 10 = 12 + 4x − 4 Now...we want to combine all the numbers with an 'x' next to them....we can do this by subtracting 4x from both sides of the equation 2x + 12 − 10 = 12 + 4x − 4 -4x -4x -2x + 12 - 10 = 12 - 4 Now we want to just combine all the other numbers -2x + 12 - 10 = 12 - 4 ( ^ ) ( ^ ) 12 - 10 is 2.......and 12 - 4 is 8 so now we have -2x + 2 = 8 Now we want to combine for the last time the numbers without the 'x' next to them so lets subtract 2 from both sides of this equation -2x + 2 = 8 -2 -2 -2x = 6 And finally ...to solve for 'x' we need to divide both sides of the equation by -2....we do this because -2x divided by -2 = 1x right? also written as just x....so -2x = 6 -------- -2 -2 x = -3 And that is your walk through :)